Verified Answer :

SJ2006

There will be 12 possible outcomes. So, the probability of (tail and a sic) would be equal to = 1/12

1/12 is your answer..........

there are 6 heads and 6 tails as and outcome so
H-T, H-T, H-T, H-T, H-T, H,T

Heads - 1

Heads - 2

Heads - 3

Heads - 4

Heads - 5

Heads - 6

Tails - 1

Tails - 2

Tails - 3

Tails - 4

Tails - 5

Tails - 6
We have to find the probability of tail and 6 so there is only one tail and a  6
P(tail and a six) = 1/12 = 0.08
hope it helps